OpenStudy (anonymous):

please help!! Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Part 2: Provide the solution to the system. 5x – 3y = 4 3x – 5y = –4

OpenStudy (anonymous):

Part 1: Line up the equations vertically. Choose one variable to eliminate, say x. Multiply the second equation by the additive inverse of the coefficient of the x of the first equation, then divide the second equation by the coefficient of the x of the second equation. Take the result and add it to the first equation (vertically) and then solve for y. Back substitute for x. Part 2: 5x – 3y = 4 3x – 5y = –4 5x – 3y = 4 -15x + 25y = 20 5x – 3y = 4 -5x + (25/3)y = 20/3 0x + (16/3)y = 32/3 16y = 32 y = 2 5x - 3(2) = 4 5x = 10 x = 2 Solution: (2,2)

OpenStudy (anonymous):

thanks so much!